# LLMs are Also Effective Embedding Models: An In-depth Overview

### v1

@@ Abstract
Large language models (LLMs) have revolutionized natural language processing by achieving state-of-the-art performance across various tasks. Recently, their effectiveness as embedding models has gained attention, marking a paradigm shift from traditional encoder-only models like ELMo and BERT to decoder-only, large-scale LLMs such as GPT, LLaMA, and Mistral. This survey provides an in-depth overview of this transition, beginning with foundational techniques before the LLM era, followed by LLM-based embedding models through two main strategies to derive embeddings from LLMs. 1) Direct prompting: We mainly discuss the prompt designs and the underlying rationale for deriving competitive embeddings. 2) Data-centric tuning: We cover extensive aspects that affect tuning an embedding model, including model architecture, training objectives, data constructions, etc. Upon the above, we also cover advanced methods, such as handling longer texts, and multilingual and cross-modal data. Furthermore, we discuss factors affecting choices of embedding models, such as performance/efficiency comparisons, dense vs sparse embeddings, pooling strategies, and scaling law. Lastly, the survey highlights the limitations and challenges in adapting LLMs for embeddings, including cross-task embedding quality, trade-offs between efficiency and accuracy, low-resource, long-context, data bias, robustness, etc. This survey serves as a valuable resource for researchers and practitioners by synthesizing current advancements, highlighting key challenges, and offering a comprehensive framework for future work aimed at enhancing the effectiveness and efficiency of LLMs as embedding models.
1 Introduction
@@
Representation learning is a key concept in deep learning, where models learn to capture meaningful features or patterns from raw data in a compressed, low-dimensional form, known as embeddings (Bengio et al., 2013; Buduma et al., 2022). In the context of information retrieval (IR), natural language processing (NLP) and computer vision (CV), representation learning is used to encode a piece of text or images into embedding vectors that capture the semantic meaning and syntactic structure of the input, enabling various downstream tasks such as classification (Muennighoff et al., 2022), retrieval (Muennighoff et al., 2022; Hao et al., 2024), clustering (Petukhova et al., 2024), recommendation (Peng et al., 2023; Liu et al., 2024b; Zhao et al., 2024), and retrieval-augmented generation (RAG) (Lewis et al., 2020; Gao et al., 2023; Caspari et al., 2024). Embeddings are crucial in modern deep learning literature because they enable efficient representation of high-dimensional data in a compact, dense format (Karpukhin et al., 2020). This compression not only reduces storage requirements but also allows for offline computation, which can then be easily used in real-time applications, including retrieval and recommendation systems, with online lightweight operations, e.g., dot-product similarity between embedded vectors. That is, embeddings preserve essential semantic and syntactic information, making it possible to perform complex operations like similarity comparison or clustering with significantly lower computational overhead.
@@
Nonetheless, pre-training Transformer encoders mainly depend on masked language modeling (MLM), where only a small proportion (e.g., 15% in BERT) of words or tokens are masked as learning objectives (Devlin et al., 2018). This deterministic learning process cannot fully perceive the rich contextual dependencies present in unmasked tokens, making these pre-training approaches less efficient and thus quickly reaching performance saturation even scaling model and data size. On the other hand, causal language modeling (CLM), learning to predict every next token given its preceding context, is more effective in utilizing both model parameters and training corpora to pre-train generative models (Radford et al., 2018), which is proven to have scaling law w.r.t. model performance. The resulting LLMs (Brown et al., 2020), mainly built upon the much deeper Transformer decoder architecture and pre-trained over trillions of tokens, have been proven to have emergent capabilities (Wei et al., 2022) in understanding and reasoning. The models, including GPT (Brown et al., 2020; Achiam et al., 2023), LLaMA (Touvron et al., 2023a; Dubey et al., 2024), and Mistral (Jiang et al., 2023a), achieve remarkable milestones in a broad spectrum of benchmarking tasks with excellent zero-/few-shot capability and state-of-the-art performance, including question-answering, coding, math, reasoning, etc.
@@
The success of LLMs has been extended to representation learning, which tunes the LLMs to generate expressively powerful embeddings (Wang et al., 2024a; Ma et al., 2024; Lee et al., 2024a; Zhuang et al., 2024; Nie et al., 2024). Intuitively, the LLMs, with significantly more parameters, larger pretraining corpora, and extended training durations, are expected to outperform the previous BERT-family models in capturing richer semantic representations and contextual nuances. This enhanced capacity enables LLMs to generate more accurate and generalizable embeddings, which is expected to improve performance across a wide range of downstream tasks compared to traditional encoder-only models.
@@
In the context of how to leverage LLMs for representation learning, there are two distinct yet complementary perspectives. On the one hand, as LLMs can be viewed as extensions of previous encoder-based models like BERT, several established techniques from the BERT era are still applicable to LLMs. For instance, methods such as corpus-aware pre-training (Li et al., 2024a), multi-task learning (Muennighoff et al., 2024), hard negative construction (Wang et al., 2024a), and distillation from cross-encoder models (Lee et al., 2024b) are still natural and practical to LLM-based approaches. Some of them have been confirmed in recent studies in terms of effectiveness, which demonstrates the utility of these techniques in fine-tuning LLMs for more robust embeddings. However, unique challenges arise when using LLMs for representation learning, particularly in determining how to extract efficient representations from a CLM (Luo et al., 2024b). Unlike BERT’s natural use of the [CLS] token to generate embeddings (Devlin et al., 2018), extracting useful representations from an LLM trained on CLM objectives (Radford et al., 2018) is less straightforward.
@@
On the other hand, the powerful expressiveness of LLMs has opened up entirely new paradigms for representation learning, one of which is termed “Direct Prompt for Embedding” (Jiang et al., 2023b; Zhuang et al., 2024). Thanks to instruction-tuning, many LLMs are endowed with the ability to follow instructions (Ouyang et al., 2022; Taori et al., 2023; Xu et al., 2023). This allows practitioners (Jiang et al., 2023b) to prompt the models to generate a specific category or a topic word, and then utilize its contextualized representations as the final embedding. This prompt-based approach presents exciting opportunities but also introduces several open research challenges. For example, the effectiveness and generalization of embeddings generated through direct prompting are still under exploration, with areas like in-context learning (Li et al., 2024b) and meta-learning (Lei et al., 2024) offering potential pathways for improving these embeddings in diverse tasks and domains.
@@
With the growing reliance on LLMs in natural language understanding and information retrieval, this survey aims to provide timely and comprehensive insights into the paradigm shift in embedding methods, offering a deeper understanding of how embedding techniques with LLMs can enhance performance across a wide range of tasks, while also highlighting critical open problems in the field. The paper begins with a brief introduction of the foundational techniques that shaped the field before the LLM era. It then introduces two primary approaches: tuning-free embedding methods, which extract meaningful text embeddings from the hidden states of LLMs through direct prompting, without the need for explicit training on embedding-specific tasks; and tuning-based embedding methods, which involve continued supervised fine-tuning, focusing on optimizing model architecture, improving training objectives, and refining training data. Next, we summarize advanced techniques developed to handle longer texts, multiple languages, and cross-modal data. We then compare the performance of various LLM-based embedding methods, considering key aspects in adapting LLMs as embedding models, such as contrasting dense versus sparse embeddings, assessing different pooling strategies, and exploring the implications of scaling laws as LLMs increase in size. Finally, the survey highlights the limitations and emerging challenges involved in adapting LLMs to be more effective as embedding models.
2 Background
@@ 2.1 Shallow Contextualization
2.2 Prominent Techniques in BERT Era
@@
Figure 1: Illustration of direct prompting or tuning LLMs as embedding models.
3 Formalization
@@
where $\mathbf{h}_{i}$ is the output hidden state or logit distribution for token $x_{i}$, a fixed-length embedding vector $\mathbf{v}$ for the entire sequence can be obtained by either taking the embedding of a special token or applying a pooling operation to the token-level embeddings.
Adapting LLMs as embedding models has become a powerful way to leverage their rich contextual understanding for various downstream tasks. This approach involves two main strategies as illustrated in Figure 1. First, LLMs can be directly prompted to generate different kinds of sequence embeddings, utilizing their inherent knowledge without additional training. Second, the quality of these embeddings can be enhanced through fine-tuning on paired or synthetic text data. These refined embeddings are versatile, supporting tasks like classification, semantic similarity, retrieval, clustering, or serving as input for other models.
4 Tuning-Free Embedding Methods
Recently, researchers have started to explore the potential of extracting meaningful text embeddings directly from the hidden states of LLMs (e.g., OPT or LLAMA) without requiring explicit training on embedding-specific tasks. These embeddings have been applied in various tasks, including clustering (Petukhova et al., 2024), recommendation (Peng et al., 2023; Liu et al., 2024b; Li et al., 2023b), and retrieval (Lei et al., 2024; Zhuang et al., 2024). These methods typically involve using fill-in-the-blank prompts. Inspired by PromptBERT, PromptEOL enhances the prompt-based method in BERT by incorporating an "explicit one word limitation" (EOL) to extract text representations for LLMs (Jiang et al., 2023b). As shown in the first line of Figure 2, PromptEOL incorporates ": “" at the end of the template to prevent the model from generating punctuation in the next token. Additionally, it considers an in-context learning framework to automatically create and search for demonstration sets to improve embeddings in LLMs. This method improves performance across all OPT models, allowing them to match or even outperform BERT in prompt-based embedding tasks.
PromptEOL (Jiang et al., 2023b)
@@ 4 Tuning-Free Embedding Methods
Figure 2: Prompts used in various tuning-free methods.
@@
More recently, Zhang et al. (2024a) revealed that EOL does not consistently yield optimal performance when fine-tuning generative models. They introduced two templates, PromptSTH and PromptSUM, which intentionally omit the “in one word” constraint, along with two powerful prompt engineering methods, Pretended CoT and Knowledge Enhancement, to enable the model to analyze diverse semantic aspects, as shown in Figure 2. To effectively capture multiple representations of sentences from distinct perspectives, Lei et al. (2024) introduced MetaEOL, which utilizes a diverse set of meta-task prompts, including text classification, sentiment analysis, paraphrase identification, and information extraction, to generate embeddings. Table 1 shows the details of the meta-tasks and their corresponding prompts. Thirukovalluru and Dhingra (2024) further presented GenEOL, which leverages LLMs to generate diverse transformations of a sentence that retain its original meaning, and then aggregates the embeddings of these transformations to enhance the final sentence embedding. Zhuang et al. (2024) directly prompted LLMs to generate both dense embedding representations and sparse bag-of-words representations for document retrieval. Springer et al. (2024) proposed echo embeddings that repeat the input twice in context and extract embeddings from the second occurrence.
Table 1: Meta-tasks and prompts utilized in PromptEOL (Zhang et al., 2024a).
Meta-Tasks Prompt.
Text Classification $\longrightarrow$ General Category Identification In this task, you’re presented with a text excerpt. Your task is to categorize the excerpt into a broad category such as ’Education’, ’Technology’, ’Health’, ’Business’, ’Environment’, ’Politics’, or ’Culture’. These categories help in organizing content for better accessibility and targeting. For this task, this sentence : "input sentence" should be classified under one general category in one word:"
Text Classification $\longrightarrow$ Opinion vs. Fact Discrimination In this task, you’re given a statement and you need to determine whether it’s presenting an ’Opinion’ or a ’Fact’. This distinction is vital for information verification, educational purposes, and content analysis. For this task, this sentence : "input sentence" discriminates between opinion and fact in one word:"
Sentiment Analysis $\longrightarrow$ Product Review Rating In this task, you’re given a review from an online platform. Your task is to generate a rating for the product based on the review on a scale of 1-5, where 1 means ’extremely negative’ and 5 means ’extremely positive’. For this task, this sentence : "input sentence" reflects the sentiment in one word:"
Sentiment Analysis $\longrightarrow$ Emotion Detection In this task, you’re reading a personal diary entry. Your task is to identify the predominant emotion expressed, such as joy, sadness, anger, fear, or love. For this task, this sentence : "input sentence" conveys the emotion in one word:"
Paraphrase Identification $\longrightarrow$ Similarity Check In this task, you’re presented with two sentences. Your task is to assess whether the sentences convey the same meaning. Use ’identical’, ’similar’, ’different’, or ’unrelated’ to describe the relationship. To enhance the performance of this task, this sentence : "input sentence" means in one word:"
Paraphrase Identification $\longrightarrow$ Contextual Synonym Detection In this task, you’re given a sentence and a phrase. Your task is to determine if the phrase can be a contextual synonym within the given sentence. Options include ’yes’, ’no’, or ’partially’. To enhance the performance of this task, this sentence : "input sentence" means in one word:"
Information Extraction $\longrightarrow$ Key Fact Identification In this task, you’re examining a news article. Your task is to extract the most critical fact from the article. For this task, this sentence : "input sentence" encapsulates the key fact in one word:"
Information Extraction $\longrightarrow$ Entity and Relation Extraction In this task, you’re reviewing a scientific abstract. Your task is to identify the main entities (e.g., proteins, diseases) and their relations (e.g., causes, treats). For this task, this sentence : "input sentence" highlights the primary entity or relation in one word:"
5 Tuning-based Embedding Model

### v2

@@ Abstract.
Large language models (LLMs) have revolutionized natural language processing by achieving state-of-the-art performance across various tasks. Recently, their effectiveness as embedding models has gained attention, marking a paradigm shift from traditional encoder-only models like ELMo and BERT to decoder-only, large-scale LLMs such as GPT, LLaMA, and Mistral. This survey provides an in-depth overview of this transition, beginning with foundational techniques before the LLM era, followed by LLM-based embedding models through two main strategies to derive embeddings from LLMs. 1) Direct prompting: We mainly discuss the prompt designs and the underlying rationale for deriving competitive embeddings. 2) Data-centric tuning: We cover extensive aspects that affect tuning an embedding model, including model architecture, training objectives, data constructions, etc. Upon the above, we also cover advanced methods for producing embeddings from longer texts, multilingual, code, cross-modal data, as well as reasoning-aware and other domain-specific scenarios. Furthermore, we discuss factors affecting choices of embedding models, such as performance/efficiency comparisons, dense vs sparse embeddings, pooling strategies, and scaling law. Lastly, the survey highlights the limitations and challenges in adapting LLMs for embeddings, including cross-task embedding quality, trade-offs between efficiency and accuracy, low-resource, long-context, data bias, robustness, etc. This survey serves as a valuable resource for researchers and practitioners by synthesizing current advancements, highlighting key challenges, and offering a comprehensive framework for future work aimed at enhancing the effectiveness and efficiency of LLMs as embedding models.
Large Language Models, Information Retrieval, Text Embedding, Prompting, Supervised Finetuning
ccs: Computing methodologies Semantic networksccs: Information systems Document representationccs: Information systems Language modelsccs: Information systems Retrieval tasks and goals
1. Introduction
@@
Representation learning is a key concept in deep learning, where models learn to capture meaningful features or patterns from raw data in a compressed, low-dimensional form, known as embeddings (Bengio et al., 2013; Buduma et al., 2022). In the context of information retrieval (IR), natural language processing (NLP) and computer vision (CV), representation learning is used to encode a piece of text or images into embedding vectors that capture the semantic meaning and syntactic structure of the input, enabling various downstream tasks such as classification (Muennighoff et al., 2022), retrieval (Muennighoff et al., 2022; Hao et al., 2024), clustering (Petukhova et al., 2024, 2025), anomaly detection (Cao et al., 2025; Gao et al., 2025), reward model (Sun et al., 2025b, a), recommendation (Peng et al., 2023; Liu et al., 2024a; Zhao et al., 2024), and retrieval-augmented generation (RAG) (Lewis et al., 2020; Gao et al., 2023; Caspari et al., 2024). Embeddings are crucial in modern deep learning literature because they enable efficient representation of high-dimensional data in a compact, dense format (Karpukhin et al., 2020). This compression not only reduces storage requirements but also allows for offline computation, which can then be easily used in real-time applications, including retrieval and recommendation systems, with online lightweight operations, e.g., dot-product similarity between embedded vectors. That is, embeddings preserve essential semantic and syntactic information, making it possible to perform complex operations like similarity comparison or clustering with significantly lower computational overhead.
@@
Nonetheless, pre-training Transformer encoders mainly depend on masked language modeling (MLM), where only a small proportion (e.g., 15% in BERT) of words or tokens are masked as learning objectives (Devlin et al., 2018). This deterministic learning process cannot fully perceive the rich contextual dependencies present in unmasked tokens, making these pre-training approaches less efficient and thus quickly reaching performance saturation even scaling model and data size. On the other hand, causal language modeling (CLM), learning to predict every next token given its preceding context, is more effective in utilizing both model parameters and training corpora to pre-train generative models (Radford et al., 2018), which is proven to have scaling law w.r.t. model performance. The resulting LLMs (Brown et al., 2020), mainly built upon the much deeper Transformer decoder architecture and pre-trained over trillions of tokens, have been proven to have emergent capabilities (Wei et al., 2022) in understanding and reasoning. The models, including GPT (Brown et al., 2020; Achiam et al., 2023), LLaMA (Touvron et al., 2023a; Dubey et al., 2024), Mistral (Jiang et al., 2023b) and Qwen (Chu et al., 2023; Zhang et al., [n. d.]), achieve remarkable milestones in a broad spectrum of benchmarking tasks with excellent zero-/few-shot capability and state-of-the-art performance, including question-answering, coding, math, reasoning, dialogue generation, etc.
@@
The success of LLMs has been extended to representation learning, which tunes the LLMs to generate expressively powerful embeddings (Wang et al., 2024c; Ma et al., 2024; Lee et al., 2024b; Zhuang et al., 2024; Nie et al., 2024). Intuitively, the LLMs, with significantly more parameters, larger pretraining corpora, and extended training durations, are expected to outperform the previous BERT-family models in capturing richer semantic representations and contextual nuances. This enhanced capacity enables LLMs to generate more accurate and generalizable embeddings, which is expected to improve performance across a wide range of downstream tasks compared to traditional encoder-only models.
@@
In the context of how to leverage LLMs for representation learning, there are two distinct yet complementary perspectives. On the one hand, as LLMs can be viewed as extensions of previous encoder-based models like BERT, several established techniques from the BERT era are still applicable to LLMs. For instance, methods such as corpus-aware pre-training (Li et al., 2024a), multi-task learning (Muennighoff et al., 2024), hard negative construction (Wang et al., 2024c), and distillation from cross-encoder models (Lee et al., 2024a) are still natural and practical to LLM-based approaches. Some of them have been confirmed in recent studies in terms of effectiveness, which demonstrates the utility of these techniques in fine-tuning LLMs for more robust embeddings. However, unique challenges arise when using LLMs for representation learning, particularly in determining how to extract efficient representations from a CLM (Luo et al., 2024b). Unlike BERT’s natural use of the [CLS] token to generate embeddings (Devlin et al., 2018), extracting useful representations from an LLM trained on CLM objectives (Radford et al., 2018) is less straightforward.
@@
On the other hand, the powerful expressiveness of LLMs has opened up entirely new paradigms for representation learning, one of which is termed “Direct Prompt for Embedding” (Jiang et al., 2023a; Zhuang et al., 2024). Thanks to instruction-tuning, many LLMs are endowed with the ability to follow instructions (Ouyang et al., 2022; Taori et al., 2023; Xu et al., 2023). This allows practitioners (Jiang et al., 2023a) to prompt the models to generate a specific category or a topic word, and then utilize its contextualized representations as the final embedding. This prompt-based approach presents exciting opportunities but also introduces several open research challenges. For example, the effectiveness and generalization of embeddings generated through direct prompting are still under exploration, with areas like in-context learning (Li et al., 2024d) and meta-learning (Lei et al., 2024) offering potential pathways for improving these embeddings in diverse tasks and domains.
@@
With the growing reliance on LLMs in natural language understanding and information retrieval, this survey aims to provide timely and comprehensive insights into the paradigm shift in embedding methods, offering a deeper understanding of how embedding techniques with LLMs can enhance performance across a wide range of tasks, while also highlighting critical open problems in the field. The paper begins with a brief introduction of the foundational techniques that shaped the field before the LLM era. It then introduces two primary approaches: tuning-free embedding methods, which extract meaningful text embeddings from the hidden states of LLMs through direct prompting, without the need for explicit training on embedding-specific tasks; and tuning-based embedding methods, which involve continued supervised fine-tuning, focusing on optimizing model architecture, improving training objectives, and refining training data. Next, we summarize advanced techniques developed to handle longer texts, multiple languages, cross-modal data, and codes. We then compare the performance of various LLM-based embedding methods, considering key aspects in adapting LLMs as embedding models, such as contrasting dense versus sparse embeddings, assessing different pooling strategies, and exploring the implications of scaling laws as LLMs increase in size. Finally, the survey highlights the limitations and emerging challenges involved in adapting LLMs to be more effective as embedding models.
2. Background
@@ 2.1. Shallow Contextualization
2.2. Prominent Techniques in BERT Era
@@
Figure 1. An overview of our survey on deploying LLMs as embedding models, covering both ① direct prompting as embedding models, ② fine-tuning as embedding models and ③ specialized embedding mdels.
3. Problem Formalization and Survey Overview
@@
$\mathbf{v}=f_{\text{agg}}(\{\mathbf{h}_{1},\dots,\mathbf{h}_{n}\}),$
where $\mathbf{h}_{i}$ is the output hidden state or logit distribution for token $x_{i}$, a fixed-length embedding vector $\mathbf{v}$ for the entire sequence can be obtained by the aggregation function $f_{\text{agg}}$, such as special-token selection (e.g., [EOS]), mean pooling, or attention-based weighting.
This formulation provides a general framework for using decoder-only LLMs as embedding models. However, practical instantiations of this process vary widely depending on how the model is used, adapted, or trained. As illustrated in Figure 1, recent research has proposed three major paradigms for generating embeddings from LLMs: direct prompting, fine-tuning, and specialization. First, we explore how LLMs can be directly prompted to produce embeddings, with a focus on prompt design and embedding derivation strategies that require no parameter updates. Second, we examine tuning-based methods that convert LLMs into dedicated embedding models through architectural adaptations (e.g., bi-attention, low-rank adaptation), refined embedding extraction mechanisms, and training objectives such as multi-task learning, multi-stage optimization, and knowledge distillation. These approaches are supported by carefully constructed datasets, leveraging both adaptation of existing data and synthesis via LLMs, often with hard negative mining to improve contrastive learning. Finally, we discuss specialized embeddings developed for multilingual, multimodal, long-context, code, reason-aware, and domain-specific scenarios, which address real-world needs and highlight the flexibility of embedding models in capturing complex semantics beyond general-purpose tasks.
In addition to these paradigms and techniques, this survey examines critical considerations in adapting LLMs as embedding models, including the comparison of dense and sparse embeddings, evaluation of various aggregation strategies, and analysis of scaling law implications. Furthermore, we also highlight open challenges and future directions, shedding light on key limitations and potential research opportunities in LLM-based embedding techniques.
4. Tuning-Free Embedding Models
PromptEOL (Jiang et al., 2023a)
@@ 4. Tuning-Free Embedding Models
Figure 2. Prompts used in various tuning-free methods.
Recently, researchers have started to explore the potential of extracting meaningful text embeddings directly from the hidden states of LLMs (e.g., OPT or LLAMA) without requiring explicit training on embedding-specific tasks. These embeddings have been applied in various tasks, including clustering (Petukhova et al., 2024), recommendation (Peng et al., 2023; Liu et al., 2024a; Li et al., 2023b), and retrieval (Lei et al., 2024; Zhuang et al., 2024). These methods typically involve using fill-in-the-blank prompts. Inspired by PromptBERT, (Jiang et al., 2023a) introduced PromptEOL, which enhances the prompt-based method in BERT by incorporating an ”explicit one word limitation” (EOL) to extract text representations for LLMs. As shown in the first line of Figure 2, PromptEOL incorporates ”: “” at the end of the template to prevent the model from generating punctuation in the next token. Additionally, it considers an in-context learning framework to automatically create and search for demonstration sets to improve embeddings in LLMs. This method improves performance across all OPT models, allowing them to match or even outperform BERT in prompt-based embedding tasks.
@@
More recently, (Zhang et al., 2024a) reveals that EOL does not consistently yield optimal performance when fine-tuning generative models. They introduced two templates, PromptSTH and PromptSUM, which intentionally omit the “in one word” constraint, along with two powerful prompt engineering methods, Pretended CoT and Knowledge Enhancement, to enable the model to analyze diverse semantic aspects, as shown in Figure 2. To effectively capture multiple representations of sentences from distinct perspectives, (Lei et al., 2024) introduced MetaEOL, which utilizes a diverse set of meta-task prompts, including text classification, sentiment analysis, paraphrase identification, and information extraction, to generate embeddings. (Thirukovalluru and Dhingra, 2024) further present GenEOL, which leverages LLMs to generate diverse transformations of a sentence that retain its original meaning, and then aggregates the embeddings of these transformations to enhance the final sentence embedding. (Zhuang et al., 2024) directly prompt LLMs to generate both dense embedding representations and sparse bag-of-words representations for document retrieval. (Springer et al., 2024) propose echo embeddings that repeat the input twice in context and extract embeddings from the second occurrence. (Li and Zhou, [n. d.]) introduced MOE Embedding (MOEE), which combines the routing weights and hidden states of Mixture-of-Experts LLMs to form a powerful zero-shot embedding model that surpasses traditional approaches without requiring any additional fine-tuning.
5. Tuning-based Embedding Models